OpenStudy (anonymous):

wdf is gbasallote 1 Medal Given x^2 - 4x +(4/2)^2 = -2 +(4/2)^2 4/2 ^2 = 4 so... x^2 - 4x + 4 =-2 + 4 x^2 - 4x +4 = 2 (x-2)^2 = 2 x-2 = sqrt 2 x = 2 plus/minus sqrt 2 wdf is sqrt 2

OpenStudy (lgbasallote):

completing the square... ax^2 + bx + c = 0 ax^2 + bx + (b/2)^2 = -c + (b/2)^2 which part did you not get?

OpenStudy (anonymous):

sqrt 2

OpenStudy (anonymous):

what does that mean

OpenStudy (lgbasallote):

(x-2)^2 = 2 you need to find x but your x is in a binomial which is squared. so first you need to get rid of it. to do that, you need to get the sqrt of (x-2)^2..which is x-2...since you square rooted the left hand side, you must do so with the right. which gives you sqrt 2

OpenStudy (anonymous):

wdf is sqrt mean?!?!?

OpenStudy (anonymous):

oh wait, square root?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

so the final answer is... ?

OpenStudy (anonymous):

the anssswer iss...?